The number of new smart phones that a phone shop sells follows a

Poisson distribution with a mean of 2 smart phones sold per day.

(i) Find the probability that the shop sells at least 4 new smart phones

during a period of 3 days. (4 marks)

(ii) Find the probability that the shop sells less than 65 new smart

phones in the month of September. (4 marks)


1
Expert's answer
2021-09-01T09:49:02-0400

Poisson distribution has the formula:

"P(X=k)=\frac{\lambda^k\cdot e^{-\lambda}}{k!}" ,

where E(X)="\lambda" =2.

We have "P(X\ge4)=1-P(X=0)-P(X=1)-P(X=2)-P(X=3)=1-e^{-\lambda}\cdot (\lambda ^0/0!+\lambda ^1/1!+\lambda ^2/2!+\lambda ^3/3!)=\\
=1-\frac{2^0/0!+2^1/1!+2^2/2!+2^3/3!}{e^2}=1-\frac{1+2+2+4/3}{e^2}=\\
=1-\frac{19}{3\cdot e^2}=0.143;"

(ii) We use sum Y of 30 random variable with Poisson distribution which has Poisson distribution also but with mean "\lambda" ="30\cdot2=60" and approximate it by normal distribution "N(60,60)" . Then "P(0\le Y\le65)=\Phi(\frac{65-60}{\sqrt{60}})+\Phi(60/\sqrt{60})=\\
\Phi(0.65)+\Phi(7.75)=0.2422+0.5=0.7422" where

"\Phi(x)" - integral function of Laplace.